Demand curve of a software development firm digisign is qd= -16p^2 -4p + 250 .Carry out analysis of demand function and find out elasticity of demand Ed At p=25.
Draw pertinent conclusions .
1
Expert's answer
2020-08-10T20:08:45-0400

Analysis

The demand function given by

"qd = -16p^2 - 4p + 250" is nonlinear. It is a quadratic demand function. The demand function is defined only in the positive quadrant.


When p = $0, "qd = -16(0^2) - 4(0) + 250"

"=250 \space units"


When qd = 0 units,

"=> 0 = -16p^2 - 4p + 250"

"=> 16p^2 + 4p - 250 = 0"

Reducing the equation by 2 gives:

"8p^2 + 2p - 125 =0"

Using the quadratic formula to find "p" gives:

"p =\dfrac { -(2) \pm \sqrt {2^2 -4(8)(-125)}} {2(8)}"


"p =\dfrac { -2 \pm \sqrt {4 004}} {16}"

"p = 3.8298230049"

"=\$3.83"


"Thus, \space qd = -16p^2 - 4p + 250 \space, \\for \space \$0 \le p \le \$3.83"


When p > $3.83, qd becomes negative. A negative quantity demanded do not make sense and hence the qd function will be undefined on price ranges that give a negative qd.


Calculating Ed

from the above analysis, the demand function is undefined when "p = \$25" because qd will be negative. Thus,

"qd = -16(25^2) - 4(25) + 250"

"= -9,850 \space units"

Thus, we can conclude that the elasticity of demand when "p = \$25" cannot be found since such point do not exist on the given demand function.